I have 98% sulfuric acid and I want a 1000ml solution of. 15% sulfuric acid. How much concentrate and how much water should I use to make my solution???
Answer by mananth(16946)

percent percentage ---------------- quantity
sulphuric acid 98 ---------------- x ml
water 0 ---------------- 1000 - x ml
Mixture 15.00% ---------------- 1000

98 x + 0 ( 1000 - x ) = 15 * 1000

98 x + 0 - 0 x = 15000
98 x - 0 x = 15000 - 0
98 x = 15000
/ 98
x= 153.06 ml 98.00% sulphuric acid
846.94 ml water